# INT-PUB-25-009, LA-UR-25-23551, PSI-PR-25-07, YITP-SB-2025-08, ZU-TH 23/25 Baryon Number Violation: From Nuclear Matrix Elements to BSM Physics

### v1

@@ 1 Introduction
Channel Limit [$10^{30}$ y] Reference
@@
Channel Limit [$10^{30}$ y] $\tau_{n\bar{n}}$ [$10^{8}$ s] Reference
@@
From the perspective of nuclear theory, these questions are intimately related to the calculation of nuclear matrix elements, though single-nucleon matrix elements or nuclear-matter effects may (also) require assessment. At the high scale, BNV can be parameterized via effective operators in SM effective field theory (SMEFT), but the observables are defined at the low, nuclear scale where the experiments are performed. For a robust calculation of these observables in terms of the underlying Wilson coefficients, one therefore needs both renormalization group evolution (RGE) and, crucially, the hadronization of the effective operators. During the workshop “INT-25-91W: Baryon Number Violation: From Nuclear Matrix Elements to BSM Physics,” held at the Institute for Nuclear Theory, University of Washington, Seattle, WA, January 13–17, 2025 [21], these interrelated questions were discussed. In particular, the workshop addressed how to extract the maximum amount of information from the available data from current and future experiments and determine their impact on the BSM landscape. In Secs. 24 we report on some of the main discussion outcomes, followed by short summaries of the contributions in Sec. 5, reflecting perspectives from experiment, EFT, lattice QCD (LQCD), and BSM model building.
@@ 2 Experiments
The workshop opened with presentations from ongoing and future experiments probing BNV including proton decay at Super-/Hyper-Kamiokande, DUNE, and JUNO as well as $n\bar{n}$ transitions at DUNE, JUNO, and HIBEAM/NNBAR. These presentations initiated detailed discussions about analysis strategies, sensitivities, and theory input for the interpretation of measurements, e.g., related to the modeling of the nuclear interaction in proton-decay searches and the conversion of $n\bar{n}$ transitions in a nucleus to the free-neutron $n\bar{n}$ transition rate. There was strong interest from the theory side to also consider channels besides the golden modes $p\to e^{+}\pi^{0}$ and $p\to\nu K^{+}$ to close other flat regions in EFT parameter space. An important revelation was that the Super-Kamiokande lifetime limits for $p\to e^{+}\gamma$ ($4.1\times 10^{34}$ y) and $p\to\mu^{+}\gamma$ ($2.1\times 10^{34}$ y) from Ref. [22] are not yet published. These limits currently do not feature in the PDG’s review [9], despite being the largest lifetime limits ever reported, exceeding the current official limits by two orders of magnitude. Accordingly, the analysis should be pushed forward to publication. In view of possible resonance enhancement of the matrix elements, the feasibility of searches with vector-meson final states was also discussed, see Sec. 3.
@@
In addition, to capture a multitude of many-body decay channels, it was proposed to conduct more inclusive searches. These inclusive searches will be less sensitive to specific channels than corresponding exclusive searches, due to the exclusive searches utilizing final-state topologic and kinetic constraints to reject atmospheric neutrino backgrounds. A proposal to report inclusive searches in energy differential bins would mitigate this disadvantage. This approach allows the experiment to recover part of the background rejection handle in kinetic constraints and facilitates easier application to specific theory-motivated channels while maintaining the nature of being inclusive. For example, in the case of $p\to e^{+}+\text{anything}$, instead of assuming two-body invisible decay [23] or certain final states (e.g., Refs. [24, 10, 25, 15]), we propose to first construct an inclusive sample with at least one $e^{+}$ in the events regardless of other features, and report the $e^{+}$ energy spectrum as well as the conversion into proton lifetime constraint assuming a mono-energetic signature. The uncertainty budget in inclusive searches is typically dominated by statistical uncertainty, due to the worse background rejection and the absence of the usually dominant uncertainty when modeling nuclear effects in proton decay. When applying the inclusive search limit to an exotic channel with $e^{+}$ in the final state that does not have a dedicated exclusive search, for example, $p\to e^{+}\pi^{0}\nu\nu$, one will first calculate the probability density spectrum of the $e^{+}$ in $e^{+}$ energy, then convolve it with the inclusive search result of $p\to e^{+}+\text{anything}$ differential in $e^{+}$ energy. Nuclear uncertainty should be applied to the probability density spectrum estimation when feasible. We envision in the future, exclusive searches will provide proton lifetime limits for theoretically well-motivated channels, while such kind of differential inclusive searches provide a wide coverage of proton decay and BNV processes.
[…]
@@ 3 Effective field theory and lattice QCD
To make the connection between the low-energy scale at which the experimental searches proceed and the high-energy scale at which the BNV physics occurs, there were two presentations featuring SMEFT. A comprehensive one-loop analysis showed that some channels were critical in resolving flat directions in parameter space, most notably the $p\to e^{+}\eta$ channel, due to different flavor structure compared to the golden mode $p\to e^{+}\pi^{0}$ [31]. For the $\eta$ channel, no direct LQCD calculations of the proton matrix elements are currently available due to the complications associated with disconnected diagrams [32]. The relevance for the SMEFT program motivates such calculations. There was also a presentation on the ongoing program to extend the RGE analysis to two-loop order [33], since the large separation of scales suggests that the sensitivity of BNV processes to two-loop corrections should be enhanced. However, it was emphasized that a consistent analysis requires the calculation of one-loop matching corrections at both the high-energy scale and in the nuclear matrix elements, i.e., even radiative corrections in the latter if QED running is to be included.
@@
In the $n\bar{n}$ context, there was a presentation highlighting the intricacies of the calculation of the nuclear conversion factor for the $n\bar{n}$ transition in deuterium, especially the sensitivity to the power counting in the nuclear EFT [34]. Currently, the precision is limited by two-nucleon contact terms, which could conceivably be determined from LQCD, as an application of the ongoing spectroscopy program in the $NN$ sector. In this way, the calculation for deuterium could be improved, the power counting tested, and better few-nucleon amplitudes provided for nuclear-structure calculations in heavier nuclei. In contrast, two- and more-nucleon effects can, to a very good approximation, be neglected for proton decay [35].
@@
Regarding proton-decay matrix elements, the feasibility of performing LQCD calculations for resonance final states such as $p\to\rho^{0}\ell^{+}$ or $p\to K^{*,+}\nu$ was discussed, given the potential extended sensitivity in case the matrix elements for these states indeed display some resonance enhancements. Critically, given the substantial theoretical effort, better projections for the expected experimental sensitivity were called for, and the $p\to K^{*,+}\nu$ channel was identified as one case in which limits could become competitive. In this regard, a potentially interesting connection to $p\to\ell^{+}\gamma$ was identified, in which vector-meson-dominance arguments could potentially be used to relate the more complicated resonance matrix elements to photon ones to get a first estimate.
@@
Finally, EFT methods can also be used to establish connections to the third generation [36], including BNV $B$-decays and proton decays involving an off-shell $\tau$. In this case, flat directions in the EFT can still be closed, but the BNV scale can be lower. On the theory side, it would be interesting to validate factorization assumptions necessary for some matrix elements with LQCD calculations, while the momentum dependence of BNV form factors can also be reconstructed using dispersion relations [37]. Moreover, in case of third-generation-induced nucleon decays, higher-multiplicity final states occur due to the decays of the $\tau$ lepton, further motivating the consideration of inclusive searches discussed in Sec. 2.
@@ 4 BSM and astrophysics connections
The landscape of possible, distinct BNV scenarios is vast, and can be organized, for instance, by $B\pm L$ and $L_{\ell}$, $\ell=e,\mu,\tau$ quantum numbers [38] ($L$ referring to total lepton number and $L_{\ell}$ that of a given flavor). Interestingly, even for higher-dimensional operators BSM realizations exist that avoid at the same time contributions at lower dimensions. To fully probe the BNV landscape, one would need to measure an unrealistically large number of channels, including high-multiplicity final states. This motivates the consideration of inclusive limits, ideally including information on the spectrum, together with a few theory-motivated exclusive modes. Besides the golden modes $p\to\pi^{0}\ell^{+}$ and $p\to K^{+}\nu$, these could include $p\to\eta\ell^{+}$ to break degeneracies in flavor space, $p\to\rho\ell^{+}$ or $p\to K^{*,+}\nu$ due to the potential resonance enhancement, and $p\to\ell^{+}\gamma$ due to large sensitivity reach. In the latter case, however, estimates of the matrix elements do suggest a suppression by the fine-structure constant $\alpha$, but refined calculations especially in view of the (unpublished) limit from Ref. [22] appear well motivated.
As specific BSM realizations, leptoquark models were discussed—in which proton decay can be absent at tree level but induced by loop effects—as were scenarios in which $n\bar{n}$ transitions would be the dominant manifestation of BNV. Finally, connections to astrophysics in the context of BNV and neutron disappearance to dark-sector final states were explored, together with possible connections to dark matter, late-scale baryogenesis models, and the tension in the neutron lifetime between beam and bottle measurements.
5 Extended abstracts
@@ 5.1 Search for proton decay in the Hyper-Kamiokande experiment
Hyper-Kamiokande is the next-generation underground water-Cherenkov detector currently under construction in the Tochibora mine, Japan, featuring a broad scientific program spanning from neutrino physics to the search for BNV. Hyper-Kamiokande is scheduled to start its operation in 2027, taking advantage of its huge fiducial mass ($186$ kton) and its new $20$" photomultiplier tubes, which feature detection efficiency and timing resolution improved by a factor $2$ with respect to Super-Kamiokande, to achieve an unprecedented sensitivity to proton decay.
@@ 5.3 Search for neutron–antineutron transition in large neutrino detectors
$n\bar{n}$ transition is a BNV process with $\Delta B=2$, providing a unique insight into the potential explanation of the baryon asymmetry in our Universe, especially in the context of post-sphaleron baryogenesis. Studies have been conducted in various neutron-rich environment including free neutron sources, neutron stars, and bound neutrons in large neutrino detectors, among which the large neutrino experiments obtained the most stringent constraints.
@@ 5.4 Search for nucleon decay in JUNO
The proton decay mode $p\rightarrow\bar{\nu}K^{+}$ is one of the two dominant decay modes predicted by a majority of GUTs [65]. In the JUNO LS detector, the decay will give rise to a three-fold coincidence feature in time, which is usually composed of a prompt signal by the energy deposit of $K^{+}$, a short-delayed signal ($\tau$ = 12.38 ns) by the energy deposit of decay daughters of $K^{+}$, and a long-delayed signal ($\tau$ = 2.2 $\mu$s) by the energy deposit of the final Michel electron. To discriminate the $p\rightarrow\bar{\nu}K^{+}$ signals from the enormous amount of backgrounds, we design a series of selection criteria based on the simulation data sample. The basic event selection about the visible energy is firstly applied. Then we employ all delayed signals selection of $p\rightarrow\bar{\nu}K^{+}$ and atmospheric neutrino events, including the Michel electron and neutron capture. Besides the common cuts on energy, position, and temporal features, additional criteria have to be explored since about 6.8% of the total atmospheric neutrino events would survive. The key part of the selections is based on the triple coincidence signature in hit time spectrum. We use the multi-pulse fitting method to reconstruct the time difference and energy of the $K^{+}$ and its decay daughters. After applying all criteria, the total efficiency for $p\rightarrow\bar{\nu}K^{+}$ is estimated to be 36.9%, while the expected background level corresponds to 0.2 events in 10 years. Thus, the JUNO sensitivity on $p\rightarrow\bar{\nu}K^{+}$ at 90% C.L. with 200 kton-years would be $\tau/\text{Br}(p\rightarrow\bar{\nu}K^{+})>9.6\times 10^{33}$ years, which is higher than the current best limit of $5.9\times 10^{33}$ yr from the Super-Kamiokande experiment [11].
We also explore the decay of bound neutrons into invisible particles (e.g., $n\rightarrow 3\nu$ or $nn\rightarrow 2\nu$) in the JUNO liquid scintillator detector, which do not produce an observable signal. The invisible decay includes two decay modes: $n\rightarrow\text{inv}$ and $nn\rightarrow\text{inv}$. The invisible decays of $s$-shell neutrons in ${}^{12}{\rm C}$ will leave a highly excited residual nucleus. Subsequently, some de-excitation modes of the excited residual nuclei can produce a time- and space-correlated triple coincidence signal in the JUNO detector. Based on a full Monte Carlo simulation informed with the latest available data, we estimate all backgrounds, including inverse beta decay events of the reactor antineutrino $\bar{\nu}_{e}$, natural radioactivity, cosmogenic isotopes, and neutral current interactions of atmospheric neutrinos. Pulse shape discrimination and multivariate analysis techniques are employed to further suppress backgrounds. With two years of exposure, JUNO is expected to give an order of magnitude improvement compared to the current best limits. After 10 years of data taking, the JUNO expected sensitivities at 90% C.L. are $\tau/\text{Br}(n\rightarrow\text{inv})>5.0\times 10^{31}\,{\rm yr}$ and $\tau/\text{Br}(nn\rightarrow\text{inv})>1.4\times 10^{32}\,{\rm yr}$.

### v2

@@ 1 Introduction
Channel Limit $10^{30}$ y] Reference
@@
Channel Limit $10^{30}$ y] $\tau_{n\bar{n}}$ $10^{8}$ s] Reference
@@
From the perspective of nuclear theory, these questions are intimately related to the calculation of nuclear matrix elements, though single-nucleon matrix elements or nuclear-matter effects may (also) require assessment. At the high scale, BNV can be parameterized via effective operators in SM effective field theory (SMEFT), but the observables are defined at the low, nuclear scale where the experiments are performed. For a robust calculation of these observables in terms of the underlying Wilson coefficients, one therefore needs both renormalization group evolution (RGE) and, crucially, the hadronization of the effective operators. During the workshop “INT-25-91W: Baryon Number Violation: From Nuclear Matrix Elements to BSM Physics,” held at the Institute for Nuclear Theory, University of Washington, Seattle, WA, January 13–17, 2025 [21], these interrelated questions were discussed. In particular, the workshop addressed how to extract the maximum amount of information from the available data from current and future experiments and determine their impact on the BSM landscape. In Secs. 24 we report on some of the main discussion outcomes, followed by short summaries of frontier developments in Sec. 5, reflecting perspectives from experiment, EFT, lattice QCD (LQCD), and BSM model building.
@@ 2 Experiments
The workshop opened with presentations from ongoing and future experiments probing BNV including proton decay at Super-/Hyper-Kamiokande, DUNE, and JUNO (see Secs. 5.1, 5.2, and 5.4), as well as $n\bar{n}$ transitions at DUNE, JUNO, and HIBEAM/NNBAR (see Secs. 5.3, 5.4, and 5.5). These presentations initiated detailed discussions about analysis strategies, sensitivities, and theory input for the interpretation of measurements, e.g., related to the modeling of the nuclear interaction in proton-decay searches and the conversion of $n\bar{n}$ transitions in a nucleus to the free-neutron $n\bar{n}$ transition rate. There was strong interest from the theory side to also consider channels besides the golden modes $p\to e^{+}\pi^{0}$ and $p\to\nu K^{+}$ to close other flat regions in EFT parameter space. An important revelation was that the Super-Kamiokande lifetime limits for $p\to e^{+}\gamma$ ($4.1\times 10^{34}$ y) and $p\to\mu^{+}\gamma$ ($2.1\times 10^{34}$ y) from Ref. [22] are not yet published. These limits currently do not feature in the PDG’s review [9], despite being the largest lifetime limits ever reported, exceeding the current official limits by two orders of magnitude. Accordingly, the analysis should be pushed forward to publication. In view of possible resonance enhancement of the matrix elements, the feasibility of searches with vector-meson final states was also discussed, see Sec. 3.
@@
In addition, to capture a multitude of many-body decay channels, it was proposed to conduct more inclusive searches. These inclusive searches will be less sensitive to specific channels than corresponding exclusive searches, due to the exclusive searches utilizing final-state topologic and kinetic constraints to reject atmospheric neutrino backgrounds. Reporting inclusive searches in energy differential bins would mitigate this disadvantage. This approach allows the experiment to recover part of the background rejection handle in kinetic constraints and facilitates easier application to specific theory-motivated channels while maintaining the nature of being inclusive. For example, in the case of $p\to e^{+}+\text{anything}$, instead of assuming two-body invisible decay [23] or certain final states (e.g., Refs. [24, 10, 25, 15]), we propose to first construct an inclusive sample with at least one $e^{+}$ in the events regardless of other features, and report the $e^{+}$ energy spectrum as well as the conversion into proton lifetime constraint assuming a mono-energetic signature. The uncertainty budget in inclusive searches is typically dominated by statistical uncertainty, due to the worse background rejection and the absence of the usually dominant uncertainty when modeling nuclear effects in proton decay. When applying the inclusive search limit to an exotic channel with $e^{+}$ in the final state that does not have a dedicated exclusive search, for example, $p\to e^{+}\pi^{0}\nu\nu$, one will first calculate the probability density spectrum of the $e^{+}$ in $e^{+}$ energy, then convolve it with the inclusive search result of $p\to e^{+}+\text{anything}$ differential in $e^{+}$ energy. Nuclear uncertainty should be applied to the probability density spectrum estimation when feasible. We envision in the future, exclusive searches will provide proton lifetime limits for theoretically well-motivated channels, while such kind of differential inclusive searches provide a wide coverage of proton decay and BNV processes.
In the context of $n\bar{n}$ and $p\bar{n}$ searches, it was observed that the inference of expected branching fractions from isospin symmetry [19] via $p\bar{p}$ [26, 27] and $n\bar{p}$ [28] experiments could be improved by decomposing the various final states into isospin components and thereby disentangling the isospin-0 and -$1$ contributions, see Ref. [29] for the $N\bar{N}\to 2\pi$ case ($N\in\{p,n\}$). The observation that $(4\text{--}6)$-pion states have the largest branching fraction can be explained by analogy to hadronic vacuum polarization $e^{+}e^{-}\to\text{hadrons}$ [30]. In this case, at $e^{+}e^{-}$ center-of-mass energies that correspond to $N\bar{N}$, multi-pion channels dominate over, e.g., the two-pion mode, since at these higher energies, resonances couple increasingly strongly to high-multiplicity channels and the threshold suppression fades away.
@@ 3 Effective field theory and lattice QCD
To make the connection between the low-energy scale at which the experimental searches proceed and the high-energy scale at which the BNV physics occurs, there were two presentations featuring SMEFT. A comprehensive one-loop analysis, see Sec. 5.6, showed that some channels were critical in resolving flat directions in parameter space, most notably the $p\to e^{+}\eta$ channel, due to different flavor structure compared to the golden mode $p\to e^{+}\pi^{0}$ [31]. For the $\eta$ channel, no direct LQCD calculations of the proton matrix elements are currently available due to the complications associated with disconnected diagrams [32]. The relevance for the SMEFT program motivates such calculations. There was also a presentation on the ongoing program to extend the RGE analysis to two-loop order [33] (Sec. 5.7), since the large separation of scales suggests that the sensitivity of BNV processes to two-loop corrections should be enhanced. However, it was emphasized that a consistent analysis requires the calculation of one-loop matching corrections at both the high-energy scale and in the nuclear matrix elements, i.e., even radiative corrections in the latter if QED running is to be included.
@@
In the $n\bar{n}$ context, there was a presentation highlighting the intricacies of the calculation of the nuclear conversion factor for the $n\bar{n}$ transition in deuterium, especially the sensitivity to the power counting in the nuclear EFT [34], see Sec. 5.8. Currently, the precision is limited by two-nucleon contact terms, which could conceivably be determined from LQCD, as an application of the ongoing spectroscopy program in the $NN$ sector, see Sec. 5.11. In this way, the calculation for deuterium could be improved, the power counting tested, and better few-nucleon amplitudes provided for nuclear-structure calculations in heavier nuclei. In contrast, two- and more-nucleon effects can, to a very good approximation, be neglected for proton decay [35].
@@
Regarding proton-decay matrix elements (see Secs. 5.10 and 5.12), the feasibility of performing LQCD calculations for resonance final states such as $p\to\rho^{0}\ell^{+}$ or $p\to K^{*,+}\nu$ was discussed, given the potential extended sensitivity in case the matrix elements for these states indeed display some resonance enhancements. Critically, given the substantial theoretical effort, better projections for the expected experimental sensitivity were called for, and the $p\to K^{*,+}\nu$ channel was identified as one case in which limits could become competitive. In this regard, a potentially interesting connection to $p\to\ell^{+}\gamma$ was identified, in which vector-meson-dominance arguments could potentially be used to relate the more complicated resonance matrix elements to photon ones to get a first estimate.
@@
Finally, EFT methods can also be used to establish connections to the third generation [36], including BNV $B$-decays and proton decays involving an off-shell $\tau$ (Sec. 5.9). In this case, flat directions in the EFT can still be closed, but the BNV scale can be lower. On the theory side, it would be interesting to validate factorization assumptions necessary for some matrix elements with LQCD calculations, while the momentum dependence of BNV form factors can also be reconstructed using dispersion relations [37]. Moreover, in case of third-generation-induced nucleon decays, higher-multiplicity final states occur due to the decays of the $\tau$ lepton, further motivating the consideration of inclusive searches discussed in Sec. 2.
@@ 4 BSM and astrophysics connections
The landscape of possible, distinct BNV scenarios is vast, and can be organized, for instance, by $B\pm L$ and $L_{\ell}$, $\ell=e,\mu,\tau$ quantum numbers [38] ($L$ referring to total lepton number and $L_{\ell}$ that of a given flavor). Interestingly, even for higher-dimensional operators BSM realizations exist that avoid at the same time contributions at lower dimensions, see Sec. 5.13. To fully probe the BNV landscape, one would need to measure an unrealistically large number of channels, including high-multiplicity final states. This motivates the consideration of inclusive limits, ideally including information on the spectrum, together with a few theory-motivated exclusive modes. Besides the golden modes $p\to\pi^{0}\ell^{+}$ and $p\to K^{+}\nu$, these could include $p\to\eta\ell^{+}$ to break degeneracies in flavor space, $p\to\rho\ell^{+}$ or $p\to K^{*,+}\nu$ due to the potential resonance enhancement, and $p\to\ell^{+}\gamma$ due to large sensitivity reach. In the latter case, however, estimates of the matrix elements do suggest a suppression by the fine-structure constant $\alpha$, but refined calculations especially in view of the (unpublished) limit from Ref. [22] appear well motivated.
As specific BSM realizations, leptoquark models were discussed (Sec. 5.14)—in which proton decay can be absent at tree level but induced by loop effects—as were scenarios in which $n\bar{n}$ transitions would be the dominant manifestation of BNV (Sec. 5.16). Finally, connections to astrophysics in the context of BNV and neutron disappearance to dark-sector final states were explored, together with possible connections to dark matter, late-scale baryogenesis models, and the tension in the neutron lifetime between beam and bottle measurements [39], see Secs. 5.15, 5.17, and 5.18.
5 Theory and experiment frontier developments
@@ 5.1 Search for proton decay in the Hyper-Kamiokande experiment
Hyper-Kamiokande is the next-generation underground water-Cherenkov detector currently under construction in the Tochibora mine, Japan, featuring a broad scientific program spanning from neutrino physics to the search for BNV. Hyper-Kamiokande is scheduled to start its operation in 2027, taking advantage of its huge fiducial mass ($186$ kton) and its new 2020" photomultiplier tubes, which feature detection efficiency and timing resolution improved by a factor $2$ with respect to Super-Kamiokande, to achieve an unprecedented sensitivity to proton decay.
@@ 5.3 Search for neutron–antineutron transition in large neutrino detectors
$n\bar{n}$ transition is a BNV process with $\Delta B=2$, providing a unique insight into the potential explanation of the baryon asymmetry in our Universe, especially in the context of post-sphaleron baryogenesis. Studies have been conducted in various neutron-rich environment including free neutron sources [20], neutron stars [58], and bound neutrons in large neutrino detectors [19, 59, 60, 61, 62, 63, 64], neutron stars, and bound neutrons in large neutrino detectors, among which the large neutrino experiments [19] obtained the most stringent constraints.
@@ 5.4 Search for nucleon decay in JUNO
[…]
We also explore the decay of bound neutrons into invisible particles (e.g., $n\rightarrow 3\nu$ or $nn\rightarrow 2\nu$) in the JUNO liquid scintillator detector, which do not produce an observable signal. The invisible decay includes two decay modes: $n\rightarrow\text{inv}$ and $nn\rightarrow\text{inv}$. The invisible decays of $s$-shell neutrons in $^{12}{\rm C}$ will leave a highly excited residual nucleus. Subsequently, some de-excitation modes of the excited residual nuclei can produce a time- and space-correlated triple coincidence signal in the JUNO detector. Based on a full Monte Carlo simulation informed with the latest available data, we estimate all backgrounds, including inverse beta decay events of the reactor antineutrino $\bar{\nu}_{e}$, natural radioactivity, cosmogenic isotopes, and neutral current interactions of atmospheric neutrinos. Pulse shape discrimination and multivariate analysis techniques are employed to further suppress backgrounds. With two years of exposure, JUNO is expected to give an order of magnitude improvement compared to the current best limits. After 10 years of data taking, the JUNO expected sensitivities at 90% C.L. are $\tau/\text{Br}(n\rightarrow\text{inv})>5.0\times 10^{31}\,{\rm yr}$ and $\tau/\text{Br}(nn\rightarrow\text{inv})>1.4\times 10^{32}\,{\rm yr}$; see Fig. 1(right).